# An effective Bombieri–Vinogradov error term for sifting problems

### v1

@@ Abstract.
In number theory, many major results related to the twin prime and Goldbach conjectures are proven using the methods of sieve theory. However, in nearly every case, the existing proofs of these results are ineffective, in that explicit values for which they hold cannot be computed. The reason for this ineffectivity is due to the reliance on the Bombieri–Vinogradov theorem. In this paper, we show that any classical sifting problem with a Bombieri–Vinogradov style error term can in fact be made effective, with no loss to the asymptotic form of the original (ineffective) result. This is done by carefully modifying the sieve upper and lower bounds as to avoid the usual complications regarding the existence of a Siegel zero. We also provide some simple applications. For example, we show that one may effectively bound the number of primes $p\leq x$ such that $p+2$ is also prime by
@@
$C_{2}=2\prod_{p>2}\left(1-\frac{1}{(p-1)^{2}}\right)$
is the twin-prime constant.
@@ 1  Introduction
As a consequence of the recent work of Sedunova [33] one can take B=A1{B=A-1} in Theorem 1.1. Up to a constant, this is the same result that one obtains upon assuming the Generalised Riemann Hypothesis (GRH), in which one has the pointwise bound (see e.g. [24, Corollary 13.8])
@@
One of the major shortcomings of the Bombieri–Vinogradov theorem is that every known proof is ineffective. In particular, it is not possible to explicitly determine how large $x$ must be in terms of $A$ to obtain a bound of the strength (1.1). Consequently, this means that any direct application of the Bombieri–Vinogradov theorem is also ineffective. The main reason for this ineffectivity is the potential existence of Siegel zero. That is, a real-valued of zero of a Dirichlet L-function very close to $1$.
In this paper we build upon recent work on developing effective variants of the Bombieri–Vinogradov theorem, which circumvent the usual complications regarding Siegel zeros. Our main results (Theorems 1.6 and 1.7) are in the context of sieve theory and thus yield effective versions of many bounds related to the twin prime and Goldbach conjectures. As a simple application, we prove the following effective upper bounds. See also Theorems 1.5 and 1.8 for examples of effective lower bound results that can be obtained via our methods.
@@
Note that heuristics for the twin prime and Goldbach conjectures suggest that (see [10, Conjectures A and B])
@@
Previously, the best effective estimates for $\Pi_{2}(x)$ and $G(n)$ gave a constant of $8$ in place of $4+\varepsilon$ in (1.4) ([34] or [29, Lemma 5]). Thus, Theorem 1.2 can be viewed as improving upon previous work by a factor of 2.
If one is not concerned with effectivity, then the constant $4+\varepsilon$ can be improved with more elaborate techniques. Most notably, in recent work by Lichtman [17, 18] and Pascadi [27], improved forms for the remainder term in the linear sieve are given, which allow one to go further than is possible with the Bombieri–Vinogradov theorem. This yields leading constants of $3.203+\varepsilon$ for $\Pi_{2}(x)$ [27, Corollary 1.4] and $3.3907+\varepsilon$ for $G(n)$ [17, Theorem 1.2]. It is likely that some variant of our results could be proven that makes Lichtman and Pascadi’s results effective, but we do not attempt to do so here.
@@ 1.1  Overview of related results
In all standard proofs of the Bombieri–Vinogradov theorem, one applies the ineffective Siegel–Walfisz theorem [24, Corollary 11.21] for small moduli $d$. However, if one is okay with a weaker bound than that in (1.1), then it is possible to apply Page’s effective bound [26, Theorem II] on Siegel zeros, in place of the Siegel–Walfisz theorem. This procedure was detailed by Liu, who obtained the following result [20, Theorem 1].
@@
Liu actually uses the logarithmic integral $\operatorname{\mathrm{li}}(x)$ in place of $\pi(x)$, but it is readily converted to the form (1.5) above by applying an effective version of the prime number theorem $\pi(x)\sim\operatorname{\mathrm{li}}(x)$ such as [15, Corollary 1.3].
@@
$\sum_{p\leq x}\left(\omega(p-1)-\log\log x\right)^{2}$
where $\omega(p-1)$ is the number of distinct prime divisors of p1p-1 with $p$ prime.
@@ 1.2  Sieve theory and the Bombieri–Vinogradov theorem
$S(\mathcal{A},\mathcal{P},z):=\left|\mathcal{A}\setminus\bigcup_{p\mid P(z)}\mathcal{A}_{p}\right|,$ (1.7)
where $p$ is prime, $z\geq 2$ and
$P(z):=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{P}\end{subarray}}p.$ (1.8)
In many settings, the Bombieri–Vinogradov theorem is used to obtain upper and lower bounds for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$. Using these upper and lower bounds, one can obtain, for example:
@@
    Upper bounds for the number of twin primes $\leq x$ [9, Theorem 3.11].
@@
    Upper bounds on the number of Goldbach representations of an even number $N$ as the sum of two primes [9, Theorem 3.11].
@@
    Upper bounds for the number of primes represented by a polynomial $F(p)$ with prime arguments [9, Theorem 5.6].
@@
    Lower bounds for the number of representations of even integers $N$ as the sum of a prime and a number with at most $2$ prime factors [9, Theorem 11.1],
@@
In the case of the application (4) above, known as Chen’s theorem, an effective and explicit version was recently obtained by Bordignon, Starickova and the author of this paper [4]. This result was obtained by significantly correcting and improving upon unpublished work of Yamada [35].
Theorem 1.5 ([4, Corollary 1.5]).
@@ 1.3  Notation and statement of main results
To state our results, we first recall the standard notation used in sieve theory. So, consider a sifting problem with a finite set of integers $\operatorname{\mathcal{A}}$ and an infinite set of primes $\operatorname{\mathcal{P}}$ such that no element of $A$ is divisible by a prime in $\overline{\operatorname{\mathcal{P}}}$. For any square-free integer $d$ with [^1] Here and throughout, by $(d,\overline{\mathcal{P}})=1$ we mean that $(d,p)=1$ for all $p\in\overline{\operatorname{\mathcal{P}}}$. $(d,\overline{\mathcal{P}})=1$, let $X>1$ (independent of $d$) and $g(d)$ be a multiplicative function with
@@
We also denote the dimension of the sifting problem as a constant $\kappa$ such that
$\prod_{\begin{subarray}{c}z_{1}\leq p<z_{2}\\ p\in\mathcal{P}\end{subarray}}\left(1-g(p)\right)^{-1}<\left(\frac{\log z_{2}}{\log z_{1}}\right)^{\kappa}\left\{1+\frac{L}{\log z_{1}}\right\}$ ($\mathbf{\Omega}(\kappa,L)$)
is satisfied for some effective constant $L>0$ and all $z_{2}>z_{1}\geq 2$. Provided $\kappa$ and $L$ are constant, it is possible to prove generic sieve upper and lower bounds of the form
@@
for some constant $\gamma_{0}>0$ with $\mu(d)$ denoting the Möbius function and $\omega(d)$ denoting the number of distinct prime factors of $d$. Examples of such sieve bounds can be found throughout any standard text such as [8] or [9]. The main principle we wish to show is that if $r(d)$ is such that the remainder sum in (1.14) is directly bounded by the Bombieri–Vinogradov theorem, then the bounds (1.12) and (1.13) can be modified to give a new remainder term $R_{D}$ of the same strength which avoids the ineffectivity due to potential Siegel zeros.
@@
Suppose that, for fixed $\kappa,L>0$, one has a generic sieve upper bound of the form (1.12) such that ($\mathbf{\Omega}(\kappa,L)$) is satisfied, with a remainder term $R_{D}$ given by (1.14) associated with a constant $\gamma_{0}>0$. Suppose then that we have a specific sifting problem $(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}})$ with constants $A\geq 1$, $\gamma_{1}>0$ such that for all $p\in\operatorname{\mathcal{P}}$,
$p\geq A+2\quad\text{and}\quad g(p)\leq A/(p-1)$ (1.15)
@@
with each implied constant effective. Let $\gamma=\gamma_{0}\gamma_{1}$. For a choice of $B>\gamma^{2}$, we set
@@
[…]
So, from Theorem 1.6, if one has a Bombieri–Vinogradov style sieve remainder term, then it is possible to get an effective upper bound for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ with little loss. In particular, as a function of $X$, the leading term in (1.18) is asymptotically the same as in (1.12). Then, in the standard case $\gamma=1$, the $O_{B,\gamma}(\cdot)$ error term in (1.18) is the same as what one would get by directly applying the Bombieri–Vinogradov theorem to $R_{D}$ in (1.12). With more work, it is likely that $B_{\gamma}$ could be improved to $B-\gamma$ for all $\gamma>0$, not just $\gamma=1$. Nevertheless, $\gamma=1$ is standard for many applications so we just use a simple but wasteful argument involving the Cauchy–Schwarz inequality for other values of $\gamma$ (see Proposition 2.12).
@@
We now move on to an effective sieve lower bound. Here, some further conditions are required. However, these conditions are still easily satisfied by most sifting problems of interest.
@@
Suppose that, for fixed $\kappa,L>0$, one has a generic sieve upper bound (1.12) and lower bound (1.13) such that ($\mathbf{\Omega}(\kappa,L)$) is satisfied, with a remainder term $R_{D}$ given by (1.14) associated with a constant $\gamma_{0}>0$. Suppose then that we have a specific sifting problem $(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}})$ with $X=|\operatorname{\mathcal{A}}|$ and such that the conditions (1.15) hold for some constant $A\geq 1$. We further suppose that for some $\operatorname{\mathcal{X}}\asymp{X\log X}$} and a constant $\gamma_{1}>0$,
@@
for all square-free $k,d>0$ with $(k,d)=1$ and $(kd,\overline{\operatorname{\mathcal{P}}})=1$. Let $\gamma=\gamma_{0}\gamma_{1}$. For some choice of $B>\max\{3,\gamma^{2}\}$, let
@@
with $\delta=o_{s,B}(1)$, every implied constant effective, and
[…]
@@
We have chosen this application as an illustrative example since its proof requires a 2-dimensional sieve and consequently has a very different setup to the proof of Theorem 1.2. Note that using a sieve-weighting procedure, the value of $17$ in Theorem 1.8 can be reduced further (see [14, Theorems 1.5 and 1.6]), but we do not do so here for brevity.
@@ 2.1  Setup and notation
Lemma 2.1 (see [24, Corollary 11.8] and [25]).
@@
Lemma 2.2 (See [7] and [28, Theorem 5.4.6]).
$\beta<1-\frac{R_{2}}{\sqrt{q}}.$
Lemma 2.3 (see [24, Corollary 11.10] or [22, Theorem 2]).
@@
For some choice of $Q\geq 2$, if the zero in Lemma 2.3 exists then we call the zero and its corresponding modulus exceptional.
Now, with a view to proving Theorems 1.6 and 1.7, we fix $\gamma>0$, $B>\gamma^{2}$, let $D(\operatorname{\mathcal{X}})=\sqrt{\operatorname{\mathcal{X}}}/(\log\operatorname{\mathcal{X}})^{B}$ and define
@@
$\mathcal{P}_{j}:=\mathcal{P}\setminus\{q_{1},\ldots,q_{j}\}\quad\text{and}\quad V_{j}(z):=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{P}_{j}\end{subarray}}\left(1-g(p)\right).$ (2.5)

### v2

@@ Abstract.
In number theory, many major results related to the additive properties of primes are proven using the methods of sieve theory. However, in nearly every case, the existing proofs of these results are ineffective, in that explicit values for which they hold cannot be computed. The reason for this ineffectivity is due to the reliance on the Bombieri–Vinogradov theorem. In this paper, we show that any classical sifting problem with a Bombieri–Vinogradov style error term can in fact be made effective, with no loss to the asymptotic form of the original (ineffective) result. This is done by carefully modifying the sieve upper and lower bounds as to avoid the usual complications regarding the existence of a Siegel zero. We also provide some simple applications. For example, we show that one may effectively bound the number of primes $p\leq x$ such that $p+2$ is also prime by
@@
$C_{2}=2\prod_{p>2}\left(1-\frac{1}{(p-1)^{2}}\right).$
This research was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship.
@@ 1  Introduction
Up to a power of $\log x$, (1.1) is the same bound as that obtained by assuming the Generalised Riemann Hypothesis, in which one has the pointwise bound (see e.g. [25, Corollary 13.8])
@@
One of the major shortcomings of the Bombieri–Vinogradov theorem is that every known proof is ineffective. In particular, it is not possible to explicitly determine how large $x$ must be in terms of $A$ to obtain a bound of the strength (1.1). Consequently, this means that any direct application of the Bombieri–Vinogradov theorem is also ineffective. The main reason for this ineffectivity is the potential existence of Siegel zeros. That is, real-valued of zeros of Dirichlet $L$-functions very close to $1$.
In this paper we build upon recent work (e.g. [1, 38, 22, 3]) on developing effective variants of the Bombieri–Vinogradov theorem, which circumvent the usual complications regarding Siegel zeros. Our main results (Theorems 1.6 and 1.7) are in the context of sieve theory and thus yield effective versions of many bounds, including those related to the twin prime and Goldbach conjectures. To make these sifting results effective, we have to carefully treat the case of “moderate” moduli (approximately $\log x\leq d\leq(\log x)^{A}$) in the Bombieri–Vinogradov theorem under the assumption of Siegel zeros. This is done by working with slightly altered sifting sets, which avoid any Siegel zeros and introduce an asymptotically small error. By contrast, in classical treatments this range of moduli is covered using Siegel’s ineffective bound for real zeros of Dirichlet $L$-functions [37].
As a simple application, we prove the following effective upper bounds. See also Theorems 1.5 and 1.8 for examples of effective lower bound results that can be obtained via our methods.
@@
Note that heuristics for the twin prime and Goldbach conjectures suggest that (see [12, Conjectures A and B])
@@
Previously, the best effective estimates for $\Pi_{2}(x)$ and $G(n)$ gave a constant of $8$ in place of $4+\varepsilon$ in (1.4) ([36] or [31, Lemma 5]). Thus, Theorem 1.2 can be viewed as improving upon previous work by a factor of 2.
If one is not concerned with effectivity, then the constant $4+\varepsilon$ can be improved with more elaborate techniques. Most notably, in recent work by Lichtman [19, 20] and Pascadi [29], improved forms for the remainder term in the linear sieve are given, which allow one to go further than is possible with the Bombieri–Vinogradov theorem. This yields leading constants of $3.203+\varepsilon$ for $\Pi_{2}(x)$ [29, Corollary 1.4] and $3.3907+\varepsilon$ for $G(n)$ [19, Theorem 1.2]. It is likely that some variant of our results could be proven that makes Lichtman and Pascadi’s results effective, but we do not attempt to do so here.
Finally, we remark that one can readily obtain strong bounds of the form (1.4) under the assumption of well-known conjectures. For example, assuming an effective form of the Elliott-Halberstam conjecture (which states that (1.1) holds for $D$ as large as $x^{1-\varepsilon}$) gives (1.4) with a leading constant of $2+\varepsilon$.
@@ 1.1  Overview of related results
In all standard proofs of the Bombieri–Vinogradov theorem, one applies the ineffective Siegel–Walfisz theorem [25, Corollary 11.21] for small moduli $d$; that is $d<(\log x)^{B}$ for any fixed $B>0$. However, if one is okay with a weaker bound than that in (1.1), then it is possible to apply Page’s effective bound [28, Theorem II] on Siegel zeros, in place of the Siegel–Walfisz theorem. This procedure was detailed by Liu, who obtained the following result [22, Theorem 1].
@@
Liu actually uses the logarithmic integral $\operatorname{\mathrm{li}}(x)$ in place of $\pi(x)$, but it is readily converted to the form (1.5) above by applying an effective version of the prime number theorem $\pi(x)\sim\operatorname{\mathrm{li}}(x)$ such as [17, Corollary 1.3].
@@
$\sum_{p\leq x}\left(\omega(p-1)-\log\log x\right)^{2}$ (1.7)
where $\omega(p-1)$ is the number of distinct prime divisors of p1p-1 with $p$ prime. Notably, (1.7) is closely related to the Titchmarsh divisor problem and a classical result of Erdős [6] on the normal order of $\omega(p-1)$.
@@ 1.2  Sieve theory and the Bombieri–Vinogradov theorem
$S(\mathcal{A},\mathcal{P},z):=\#\{a\in\operatorname{\mathcal{A}}:p\in\operatorname{\mathcal{P}}\ \text{and}\ p\mid a\implies p\geq z\}.$ (1.8)
In many settings, the Bombieri–Vinogradov theorem is used to obtain upper and lower bounds for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$. As a (non-exhaustive) list of examples, one can use these upper and lower bounds to obtain:
@@
    Upper bounds for the number of twin primes $\leq x$ [11, Theorem 3.11].
@@
    Upper bounds on the number of Goldbach representations of an even number $N$ as the sum of two primes [11, Theorem 3.11].
@@
    Upper bounds for the number of primes represented by a polynomial $F(p)$ with prime arguments [11, Theorem 5.6].
@@
    Lower bounds for the number of representations of even integers $N$ as the sum of a prime and a number with at most $2$ prime factors [11, Theorem 11.1],
@@
In the case of the application (4) above, known as Chen’s theorem, an effective and explicit version was recently obtained by Bordignon, Starichkova and the author of this paper [3]. This result was obtained by significantly correcting and improving upon unpublished work of Yamada [38].
Theorem 1.5 ([3, Corollary 1.5]).
@@ 1.3  Notation and statement of main results
To state our results, we first recall some standard notation used in sieve theory. So, consider a sifting problem with a finite set of integers $\operatorname{\mathcal{A}}$ and an infinite set of primes $\operatorname{\mathcal{P}}$ such that no element of $\operatorname{\mathcal{A}}$ is divisible by a prime in $\overline{\operatorname{\mathcal{P}}}=\{p\ \text{prime}:p\notin\operatorname{\mathcal{P}}\}$. For any square-free integer $d$ with [^1] Here and throughout, by $(d,\overline{\mathcal{P}})=1$ we mean that $(d,p)=1$ for all $p\in\overline{\operatorname{\mathcal{P}}}$. $(d,\overline{\mathcal{P}})=1$, let $X>1$ (independent of $d$) and $g(d)$ be a multiplicative function with
@@
We also denote the dimension of the sifting problem as a constant $\kappa>0$ such that
$\prod_{\begin{subarray}{c}z_{1}\leq p<z_{2}\\ p\in\mathcal{P}\end{subarray}}\left(1-g(p)\right)^{-1}<\left(\frac{\log z_{2}}{\log z_{1}}\right)^{\kappa}\left\{1+\frac{L}{\log z_{1}}\right\}$ (1.12)
is satisfied for some effective constant $L>0$ and all $z_{2}>z_{1}\geq 2$. Provided $\kappa$ and $L$ are constant, one can prove generic sieve upper and lower bounds of the form
@@
for some constant $\gamma_{0}>0$ with $\mu(d)$ denoting the Möbius function and $\omega(d)$ denoting the number of distinct prime factors of $d$. Examples of such upper and lower bounds can be found throughout any standard text on sieve theory including [11, 10] or [8] and are referred to as variants of “the fundamental lemma of sieve theory”. The main principle we wish to show is that if $r(d)$ is such that the remainder sum in (1.15) is directly bounded by the Bombieri–Vinogradov theorem, then the bounds (1.13) and (1.14) can be modified to give a new remainder term $R_{D}$ of the same strength which avoids the ineffectivity due to potential Siegel zeros.
@@
Suppose that, for fixed sieving dimension $\kappa>0$ one has a generic sieve upper bound of the form (1.13) with a remainder term $R_{D}$ given by (1.15) associated with a constant $\gamma_{0}>0$. Suppose then that we have a specific sifting problem $(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}})$ with constants $A\geq 1$, $\gamma_{1}>0$ such that for all $p\in\operatorname{\mathcal{P}}$,
$g(p)\leq A/(p-1)$ (1.16)
@@
with each implied constant effective. Let $\gamma=\gamma_{0}\gamma_{1}$. For a choice of $B>\gamma^{2}+1$, let
@@
$B_{\gamma}=\begin{cases}B-2,&\text{if}\ 0<\gamma\leq 1,\\ \frac{B-1-\gamma^{2}}{2},&\text{if}\ \gamma>1.\end{cases}$ (1.20)
So, from Theorem 1.6, if one has a Bombieri–Vinogradov style sieve remainder term, then it is possible to get an effective upper bound for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ with little loss. In particular, as a function of $X$, the leading term in (1.19) is asymptotically the same as in (1.13). Then, in the standard case $\gamma=1$, the $O_{B,\gamma}(\cdot)$ error term in (1.19) is the same as what one would get by directly applying the Bombieri–Vinogradov theorem to $R_{D}$ in (1.13). With more work, it is likely that $B_{\gamma}$ could be improved to $B-1-\gamma$ for all $\gamma>0$, not just $\gamma=1$. Nevertheless, $\gamma=1$ is standard for many applications so we just use a simple but wasteful argument involving the Cauchy–Schwarz inequality for other values of $\gamma$ (see Proposition 2.11).
@@
We now move on to an effective sieve lower bound. Here, the conditions required are slightly more restrictive. However, these conditions are still easily satisfied by most sifting problems of interest.
@@
Suppose that one has a generic sieve upper bound (1.13) and lower bound (1.14) such that (1.12) is satisfied for a fixed $\kappa,L>0$, with a remainder term $R_{D}$ given by (1.15) associated with a constant $\gamma_{0}>0$. Suppose then that we have a specific sifting problem $(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}})$ with $X=|\operatorname{\mathcal{A}}|$ and such that the condition (1.16) holds for some constant $A\geq 1$. We further suppose that for some $\operatorname{\mathcal{X}}\asymp{X\log X}$} and a constant $\gamma_{1}>0$,
@@
for all square-free $k,d>0$ with $(k,d)=1$ and $(kd,\overline{\operatorname{\mathcal{P}}})=1$. Let $\gamma=\gamma_{0}\gamma_{1}$. For some choice of $B>\gamma^{2}+1$, let
@@
with $\delta=o_{s,B}(1)$, every implied constant effective, and $B_{\gamma}$ is as in (1.20).
@@
We have chosen this application as an illustrative example since its proof requires a 2-dimensional sieve and consequently has a very different setup to the proof of Theorem 1.2. Using a sieve-weighting procedure, the value of $17$ in Theorem 1.8 can be reduced further (see [16, Theorems 1.5 and 1.6]), but we do not do so here for brevity.
@@ 2.1  Setup and notation
For any fixed $q$, the function $\mathcal{L}_{q}(s)$ can have at most $1$ “bad” real zero which is close enough to the $1$-line as to significantly worsen estimates on the number of primes in an arithmetic progression mod $q$. This is made precise by the following standard result.
Lemma 2.1 (see [25, Corollary 11.8] and [27]).
@@
Another classical result related to Lemma 2.1, is that amongst all moduli $q$ up to some height $Q$, there is at most one “really bad” Siegel zero. This is useful for our purposes since the Bombieri–Vinogradov theorem is an averaging statment over a range of moduli.
Lemma 2.2 (see [28, Lemma 8] or [24, Theorem 2]).
@@
For some choice of $Q\geq 2$, if the zero in Lemma 2.2 exists then we call the zero as well as its corresponding modulus and character exceptional.
The final preliminary result that we require is an effective bound showing that Siegel or exceptional zeros cannot be “too close” to the 1-line.
Lemma 2.3 (See [9] and [30, Theorem 5.4.6]).
$\beta<1-\frac{R_{2}}{\sqrt{q}}.$ (2.2)
Remark.
Lemma 2.3 is a strengthening of Page’s theorem [28, Theorem II], which gives $1-\beta=O(1/\sqrt{q}(\log q)^{2})$. Although Page’s weaker bound would be sufficient for our purposes, the log-free bound (2.2) is neater and simpler to work with.
Now, with a view to proving Theorems 1.6 and 1.7, we fix $\gamma>0$, $B>\gamma^{2}+1$, let $D(\operatorname{\mathcal{X}})=\sqrt{\operatorname{\mathcal{X}}}/(\log\operatorname{\mathcal{X}})^{B}$ and define
@@
$\mathcal{P}_{j}:=\mathcal{P}\setminus\{q_{1},\ldots,q_{j}\}\quad\text{and}\quad V_{j}(z):=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{P}_{j}\end{subarray}}\left(1-g(p)\right)$ (2.6)